change the following fraction into decimal, write the solution. 1. 3/8 , 2.6/12 3. 4/12 4. 9/16 5. 8/26
Answer by CubeyThePenguin(3113)

3/8 = 0.375
6/12 = 0.5
4/12 = 0.333...
9/16 = 0.5625
8/26 = 0.307692...